OpenStudy (anonymous):

what is the length of the leg of a right triangle are 25 cm and 60 cm.what is the length of the hypotenuse?

OpenStudy (anonymous):

a^2 + b^2 = c^2 (c is always the hypotenuse)....solve for c

OpenStudy (anonymous):

is it 65 cm 35 cm 42.cm or 9.2cm if you coud tell me i need the awnser for a online test and its the last question and i have 3 min

OpenStudy (anonymous):

25^2 + 60^2 is 625 + 3600 = c^2 c = sqrt(4225)

OpenStudy (anonymous):

thank you gator the answer is 65 cm